OpenStudy (anonymous):

z=2/3(x^3/2+y^3/2), 0 less than or equal to x less than or equal to 5, 0 less than or equal to y less than or equal to 3, I worked it out as a double integral and got 65.5 but does not match any of the answer choices on my sheet was wondering if someone one could work it out and see what they get. Thanks

ganeshie8 (ganeshie8):

im getting 190

OpenStudy (anonymous):

answer choices are as follow

OpenStudy (anonymous):

36.238 38.031 31.017 29.973 33.018

OpenStudy (anonymous):

not sure what i am doing wrong no instruction or directions are given in the problem just that

OpenStudy (anonymous):

not sure what else to do just though it would be a double integral and that is all

OpenStudy (anonymous):

hello?

ganeshie8 (ganeshie8):

may i knw why are u evaluating doule integral... and whats the complete quesiton ?

OpenStudy (anonymous):

friend told me to do it that way and I tried it now that I got home

OpenStudy (anonymous):

that is the entire question

ganeshie8 (ganeshie8):

could u possibly take a snapshot and attach ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ill have it in a bit hopefully if i learn the hang of this chromebook lol

OpenStudy (anonymous):

is that ok enough for you

OpenStudy (anonymous):

ITs a bonus question and when I printed it the top with the instructions was cut off

OpenStudy (anonymous):

ill ask him real quick what it says

ganeshie8 (ganeshie8):

oh ok, im also getting the same answer as u : http://www.wolframalpha.com/input/?i=%5Cint_0%5E5+%5Cint_0%5E3++2%2F3%28x%5E%283%2F2%29%2By%5E%283%2F2%29%29+dy+dx

OpenStudy (anonymous):

professor said he would make the five bonus problems calculus 3 problem even though we are in calculus 2 I thought you learned double integrals thast why my friend told me that yea I should try it that way but let me ask this other guy if the instructions printed out

OpenStudy (anonymous):

okay he said that it says something about area of surface

OpenStudy (anonymous):

hope you undertood that english is not my first language so sorry

OpenStudy (anonymous):

not sure what it means by area of surface havent learned that yet hope to learn it next year though but if you know how to do it I would like to get a preview right now if thats ok with you

OpenStudy (anonymous):

he sent me this hope that clarifys it a bit its not of high quality but it is sort of readable

ganeshie8 (ganeshie8):

yes it clarifies everything :)

OpenStudy (anonymous):

this is next level stuff so I do not ndertsand it but if you could kinda of explain it to me to prepare for next year( thats what he wants us to do the professor) i would be happy about thhat

ganeshie8 (ganeshie8):

sure :) first let me just say what u were trying earlier : 65.5 is the VOLUME under the given function. Its not the surface area.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

would you do something like triple integrating to fine the surface area or what would you do

ganeshie8 (ganeshie8):

nope, double integral will do

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

is there an extra step I have to take?

ganeshie8 (ganeshie8):

do u knw how to find "arc length" in calcII ?

OpenStudy (anonymous):

I honestly do not remember covering that but let me take a look at my notes

OpenStudy (anonymous):

theres nothing here on arc length sorry

ganeshie8 (ganeshie8):

no need, its not necessary

OpenStudy (anonymous):

we have done integration l hopitals series and other stuff like that

OpenStudy (anonymous):

also some basic diff equations

OpenStudy (anonymous):

but no arc lenght

OpenStudy (anonymous):

oh so its not needed then

OpenStudy (anonymous):

so how would we do it? sorry

OpenStudy (anonymous):

am i still talking with someone I have no clue how this website works its my first time sorry

ganeshie8 (ganeshie8):

look up the link i gave u earlier

OpenStudy (anonymous):

the video link

OpenStudy (anonymous):

that one?

ganeshie8 (ganeshie8):

if you want the answer : 33.018 http://www.wolframalpha.com/input/?i=%5Cint_0%5E3+%5Cint_0%5E5+sqrt%28x%2By+%2B+1%29+dx+dy

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

For answer and vidoe link

OpenStudy (anonymous):

one more question though for a volume problem like this, find the volume under z=x^5+y^5 and above the region bounded by y=x^2 and x=y^2 would you do just a double integral then

OpenStudy (anonymous):

you said that when i did the double integral for the last problem and got 65 i believe it was to get the voluem

OpenStudy (anonymous):

would I do the same here for that problem

OpenStudy (anonymous):

and how would I set it up if it were a double integral problem I am jsut beginning with these double integrala nd would not knwo how to set this one up

ganeshie8 (ganeshie8):

yes but u need to setup the bounds correctly

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

but how would you set it up im sorry

OpenStudy (anonymous):

wow I am already seeing how hard cal 3 is going to be its start to scare me hehe

ganeshie8 (ganeshie8):

its not hard if u start from beginning

OpenStudy (anonymous):

became a fan of you hehe you are very helpful

OpenStudy (anonymous):

really its not hard

ganeshie8 (ganeshie8):

i like that video ^ it teaches EVERYTHING about double integrals in a simple manner in just 30 minutes

ganeshie8 (ganeshie8):

watch it and let me knw if u have any questions

OpenStudy (anonymous):

i will thanks

OpenStudy (anonymous):

but can you jsut fnish off these problem with me I have to go to bed after this and work tomorrow hehe

ganeshie8 (ganeshie8):

we're already done wid the current problem right ?

OpenStudy (anonymous):

i meant fir the new problem though

ganeshie8 (ganeshie8):

whats the new problem ?

OpenStudy (anonymous):

it was find the volume under z= x^5+y^5 and above the region bounded by y=x^2 and x=y^2

OpenStudy (anonymous):

did I send it , i might have not

OpenStudy (anonymous):

maybe i just thought i did ...sorry if i did

OpenStudy (anonymous):

I was talking about this one for the last couple of minutes

OpenStudy (anonymous):

weird thats not an aswer choice

OpenStudy (anonymous):

the choices are a.1/37 b.1/18 c.1/32 d.3/32 e.5/32

ganeshie8 (ganeshie8):

wat are the options ?

OpenStudy (anonymous):

the answer you gave me is 5/32

OpenStudy (anonymous):

thats not equal to any of the fractions given above is it?

OpenStudy (anonymous):

let me grab my calculator and check

OpenStudy (anonymous):

1/18 is the closest

ganeshie8 (ganeshie8):

not sure 3/52 has to be the answer

ganeshie8 (ganeshie8):

check the question once

OpenStudy (anonymous):

it says

OpenStudy (anonymous):

Find the volume under z=x^5+y^5 and above the region bounded by y=x^2 and x=y^2

OpenStudy (anonymous):

it matches what I wrote previously

OpenStudy (anonymous):

might have been a typo on the answer choices

ganeshie8 (ganeshie8):

okay 3/52 is the correct answer, and its not there in ur options.

OpenStudy (anonymous):

im just gonna make a note of it

OpenStudy (anonymous):

for the professor

OpenStudy (anonymous):

thanks again for your help and are you here tomorrow as well

OpenStudy (anonymous):

would like it if you can help me on the remaining problems I have

OpenStudy (anonymous):

and what did you study in college that you know so much hehe?

ganeshie8 (ganeshie8):

il be around, but id like u to watch the videos first

ganeshie8 (ganeshie8):

after calcII, i just watched below videos one by one : http://ocw.mit.edu/courses/mathematics/18-02-multivariable-calculus-fall-2007/video-lectures/

OpenStudy (anonymous):

Yes I am going to try to watch the videos for sure if I can get around to it tomorrow after work

OpenStudy (anonymous):

if not could you still help me this work is due monday for cal 2 hehe